# PANTS: Practical Adversarial Network Traffic Samples against ML-powered Networking Classifiers

### v1

@@ Abstract
Multiple network management tasks, from resource allocation to intrusion detection, rely on some form of ML-based network-traffic classification (MNC). Despite their potential, MNCs are vulnerable to adversarial inputs, which can lead to outages, poor decision-making, and security violations, among other issues.
@@
To address these challenges, we introduce PANTS, a practical white-box framework that uniquely integrates AML techniques with Satisfiability Modulo Theories (SMT) solvers to generate adversarial inputs for MNCs. We also embed PANTS into an iterative adversarial training process that enhances the robustness of MNCs against adversarial inputs. PANTS is 70% and 2x more likely in median to find adversarial inputs against target MNCs compared to two state-of-the-art baselines, namely Amoeba and BAP. Integrating PANTS into the adversarial training process enhances the robustness of the target MNCs by 52.7% without sacrificing their accuracy. Critically, these PANTS-robustified MNCs are more robust than their vanilla counterparts against distinct attack-generation methodologies.
@@ 1 Introduction
Managing networks involves tasks such as performance routing, load balancing, and intrusion detection and is extremely challenging due to the highly-diverse, heavy-tailed, and unpredictable inputs, i.e., network traffic. ML-based network-traffic classification (MNCs) provides a compelling alternative for handling these complex tasks more efficiently [18, 33, 11, 32, 23, 10, 20, 19, 43]. However, like other ML-based solutions [35, 15], MNCs are vulnerable to adversarial inputs, that is, inputs meticiously perturbed to deceive the MNC. Attackers can exploit this to compromise critical network infrastructure by subtly perturbing traffic characteristics like packet sizes and inter-arrival times, leading to security breaches or downtime. Beyond being dangerous, these attacks are particularly practical in the networking domain, where multiple entities have access to network traffic.
@@
What is critically lacking is a systematic framework for network operators and MNC designers to enhance the robustness of MNCs against attackers. This involves identifying adversarial inputs, meaning perturbations on original packet sequences that deceive the target MNC into misclassifying them. Operators can then use these inputs to: (i) refine and optimize the MNC by exploring different models, architectures, or features; (ii) assess the attack surface of their MNCs, understand the potential consequences and plan accordingly; and (iii) use them to strengthen MNCs against realistic threat models through re-training or fine-tuning. Critically, operators will not use or take the framework seriously if the generated adversarial inputs are (i) not realizable, meaning they do not fall within a realistic attacker’s capabilities, or (ii) do not retain the semantics of the original sequence, meaning the ground-truth label of the perturbed inputs has changed. In fact, we show that realizable and semantics-preserving adversarial inputs are more useful for adversarial training compared to purely adversarial ones.
@@
While there are numerous tools for generating adversarial samples in the Adversarial Machine Learning (AML) literature [16, 27, 7] they are not directly applicable to MNCs. First, AML tools are built for systems that are end-to-end differentiable, which makes them unsuitable for typical MNCs that include a feature engineering module to transform packet sequences into feature vectors before they are fed to the ML model. Although AML tools can generate adversarial feature vectors, there is no guarantee these vectors can be reverse-engineered into valid packet sequences. Even if a valid packet sequence exists, the perturbations needed to convert the original sequence into an adversarial one may exceed what a realistic attacker could achieve. Second, it is uncertain whether AML-generated samples are preserving the semantics of the original packet sequence while deceiving the MNC. Although AML techniques often incorporate a perturbation budget (expressed as a norm), this concept does not translate well to networking traffic. For example, perturbing a VoIP flow (packet sequence) to resemble web browsing by delaying packets might deceive a traffic classifier but degrades the original flow’s performance, making the call unusable. Critically, adversarially training an MNC with non-realizable or non-semantics-preserving samples (i.e., directly using what AML generates) will degrade the MNC’s accuracy without significantly improving robustness as we show in §2.
@@
To address these shortcomings, this paper presents PANTS, a framework to generate Practical Adversarial Network Traffic Samples that are realizable for various threat models and semantic-preserving. PANTS integrates traditional methods for generating adversarial samples (AML), such as Projected Gradient Descent (PGD) [27] or Zeroth-order Optimization (ZOO) [7], with formal methods, concretely a Satisfiability Modulo Theories (SMT) solver. The AML perturbs the original inputs (packet sequences or features) in a manner that maximizes their distance (e.g., loss for MLP) from the original decision boundaries with consistently high success rate. The SMT solver finds packet sequences that do not diverge too much from the AML’s output but are realizable, and consistent with the threat model and preserve the semantics of the original packet sequence. To achieve this, we encode the feature engineering, the threat model, and other networking constraints needed for semantic preservation into logical formulas. Among other optimizations, PANTS iteratively adds and removes constraints that come from the AML, prioritizing those that have a substantial effect on confusing the classifier. Finally, we integrate PANTS into an interactive training process that enhances the target MNC.
We have evaluated PANTS against three distinct ML-powered network classifiers, each implemented with different models and processing pipelines. PANTS finds adversarial, realizable and semantic-preserving samples with 35.31% success rate in median, that is 70% and 2x higher than two SOTA methods. Despite being a white-box approach, PANTS works well with gradient (e.g., transformer encoder) and non-gradient (e.g., random forest) models and various processing pipelines, including those containing non-differentiable components (e.g., feature engineering). Importantly, we integrate PANTS into a novel adversarial training pipeline, which improves MNCs’s robustness against both PANTS-generated inputs and Amoeba-generated inputs. We will, of course, make PANTS and our use cases publicly available to facilitate further research and to improve reliance on MNCs.
Ethics Statement: This work doesn’t raise any ethical issues.
@@ 2 Motivation & Limitations of Existing Work
Figure 1: MNCs are vulnerable to adversarial inputs (circles). Traditional adversarial training sacrifices accuracy to reduce vulnerability (star). Our iterative adversarial-training process that leverages PANTS reduces the vulnerability of MNCs without hurting their accuracy (cross).
@@ 2.1 Motivating use case
Consider a network operator that uses an ML-based traffic classifier to distinguish real-time applications (e.g., online gaming). The operator uses this information to optimize routing: real-time applications use fast but low-bandwidth network paths. This is one of the multiple network management tasks for which network operators could use a ML-based traffic classifier (built with Multilayer Perceptron-MLP, Random Forest, Transformers [25, 6, 18, 33, 11]) and benefit from its ability to learn complex patterns directly from data. While highly useful, this MNC exposes the network to various exploits. Two representative examples include attackers controlling (i) the end hosts of the communication or (ii) an in-network component. In the former case, a malicious (or simply “rational”) application developer could manipulate the traffic generated by her application to be misclassified as gaming by the network’s MNC, hence gaining an unfair performance advantage over competitors. This could be achieved by subtly altering the size of some packets or delaying specific packets without significantly impacting the overall performance of her application’s traffic. In the latter case, a malicious or compromised router in-path can increase the volume of traffic routed through the fast but low-bandwidth path by subtly altering the non-real-time traffic to cause the MNC to misclassify it as real-time. These perturbations would result in congestion in the low-bandwidth path, leading to delays or even downtime for actual real-time traffic, which is vital for the network’s profitability. The attacker cannot inject packets because that would be detectable from the application perspective, but she can delay packets.
ML network classifiers are vulnerable to adversarial inputs. Fig. 1 illustrates the likelihood of an attacker controlling end hosts (e.g., a malicious app developer) to successfully manipulate their traffic to cause an MNC to misclassify it (predict a favorable for the attacker label). The results show that, despite the high accuracy of the models during testing, the attacker can deceive classifiers implemented with MLP, CNN, and a transformer with success rates of 81.12%, 94.25% and 99.80%, respectively. While we identified these manipulations (adversarial inputs), hence the vulnerability of the MNCs using PANTS, which is a white-box approach with complete access to the MNC’s implementation, this does not diminish the significance of the risk they pose. Security through obscurity is never a sound strategy, and network operators should be prepared for (or at least aware of) the worst-case scenario where an attacker has complete knowledge of their system.
@@ 2.2 Requirements
Adversariness & Semantic preservation .  The framework needs to generate inputs that are adversarial, meaning that they are capable of misleading MNCs into producing an incorrect classification. To ensure that the classification is indeed incorrect, these inputs should be created by meticulously applying perturbations to the original labeled samples without altering their semantics. In other words, each adversarial input should still represent the same underlying meaning or class as the original, which is described as a set of constraints on the adversarial input. The closest notion of semantics in the image domain is the perturbation budget, which sets a maximum allowable distance between the original and perturbed image [27]. However, such a metric doesn’t translate directly to the networking domain, where there are dependencies across packets or features of each input, which perturbations should respect. We assume that semantics can be expressed as constraints on the variables of the input (not only as a distance between the adversarial and the original), and the framework should provide the means to respect them
@@ 2.3 Limitations of existing work
Synthetic data generation Networking-specific generative models such as NetShare [44] and Netdiffusion [21] have shown strong promise in generating high-fidelity synthetic traffic. While synthetic traces can, in some cases, be used to improve the performance of ML models [22], the synthetic traces are unlikely to mislead an ML classifier, and even if they do, one has no way of knowing the correct label.
Black-box RL-based techniques While useful for aspiring attackers, black-box RL-based approaches [14] for generating adversarial examples against MNCs are incapable of meeting our requirements. Amoeba [26], for example, trains an agent to perturb each packet to maximize the likelihood that the altered flow will deceive the MNC. Amoeba fulfills the generality requirement as it is black-box, and can fulfill realizability if trained with the right actions. Still, Amoeba’s outputs are not always semantic-preserving and adversarial because there is no way to guarantee those during inference. Also, as we show in our evaluation, Amoeba’s generated adversarial inputs are less effective in robustifying MNCs during adversarial training, possibly because it (over)exploits a vulnerability rather than trying to find new ones. This is a reasonable strategy for an attacker, which is the focus of such a black-box approach.
White-box gradient-based techniques Among the group of white-box and gradient-based approaches (which includes AML that we discuss in §3.3), a representative example (due to the focus on networking inputs) is BAP [29]. BAP adversarially trains a generator to create perturbation (e.g., delay, number of added dummy bytes) on every single packet of a packet sequence such that the distance (e.g., loss) between the predicted label of the perturbed sequence and its ground truth is maximized. Because BAP uses gradient for maximizing the loss it does not work for non-end-to-end differentiable MNCs, hence failing the generality requirement. Further, by perturbating each packet independently, we cannot enforce semantics, which can be flow-wide. We compare against BAP in §5.
@@ 3 Overview
Figure 2: Overview of PANTS workflow. PANTS generates adversarial inputs that are also used to iteratively fine-tune the target MNC. PANTS receives the implementation of a MNC together with a training dataset and a couple of rules that constrain the generated inputs. At its core, PANTS features an AML component that collaborates with an SMT solver.
@@ 3.1 Problem formulation
We define a threat model as a set of pertubations, $\delta$ ($\delta:P\rightarrow P$) on the original packet sequences, $p_{o}$, to generate adversarial packet sequences $p_{a}$, ($p_{a}=\delta(p_{o})$). $\delta(p_{o})$ is adversarial if $f(\phi(\delta(p_{o})))\neq f(\phi(p_{o}))$, meaning that the perturbed flow will be misclassified by $f$ and $\delta(p_{o})$ preserves the semantics of $p_{o}$. Semantics are flow-level constraints on $x_{o}$ i.e., constraints that can be dependent on the whole packet sequence not just each packet indecently.
Our goal is to design a framework that generates packet sequences $p_{a}$ from $p_{o}$ such that $p_{a}$ are adversarial to $f$, semantic-preserving to $p_{o}$, and are within the capabilities of a threat model.
Figure 3: Given a network flow p, the perturbation $\delta$ is applied to generate an adversarial flow $\delta(f)$, which can cause the ML model to return a wrong output.
@@ 3.2 The promise of Adversarial ML
Figure 4: PANTS combines AML with SMT to generate adversarial and realizable samples (flows).
3.3 AML limitations
@@
However, in the networking domain, the traffic is often processed by the feature engineering module, which generates feature vectors that are fed to the ML model. While gradient information can still guide the modification of each feature to increase loss, it is unclear how to map modifications to the feature vector into the corresponding packet sequence as the feature engineering is non-differential and non-invertible. In fact, there is no guarantee that a perturbed feature vector would have a corresponding valid packet sequence as AML changes each feature independently, making contradicting decisions. For example, suppose an ML-powered intrusion detection system whose feature set includes both minimum and maximum packet inter-arrival times ($min\_iats$, $max\_iats$). Since AML does not understand the two features are dependent, in other words, that $min\_iat\leq max\_iats$ must always hold, it is possible for AML to decrease the value of $max\_iats$ and increase $min\_iats$ such that $min\_iats>max\_iats$. In this case, while the generated feature vector might be adversarial in that it might be misclassified by the intrusion detection system, it is not a realizable sample i.e., there is no packet sequence that would ever result in this feature set. In short, using the perturbed feature vectors is like defending against an impossible attack. While one can try to use a mask [30] to only allow AML to change a subset of the features that are not dependent, that would reduce the inputs that AML can find. Indeed, while AML can run with as few as a single feature there is no guarantee that the generated sample would be adversarial.
@@
Even if the target model does not use (non-differentiable) feature engineering or if one somehow finds a packet sequence that is consistent with the AML adversarial feature vector, the packet sequence might still fall outside the capabilities of the threat model, violate any network constraints, or break the semantics of the original flow. For example, suppose the threat model only allows injecting packets, and flow duration $f_{1}$ and packets per second $f_{2}$ are two features extracted by the feature engineering module. The number of packets can be derived from #$\#packets=f_{1}/f_{2}$. If #$\#packets$ derived from the adversarial features is less than the number of packets in the original flow, even though one can find an adversarial packet sequence consistent with the adversarial features, it still falls outside the capabilities of the threat model. Similarly, for a deep NN that is directly accepting the packet sequence as input, the AML-generated adversarial sequence could change the interarrivals in a way that is impossible for any congestion control algorithm or in a way that makes the connection unusable because it is too slow in the beginning.
3.4 PANTS: AML for networking

### v3

# Robustifying ML-powered Network Classifiers with PANTS

@@ Abstract
Multiple network management tasks, from resource allocation to intrusion detection, rely on some form of ML-based network traffic classification (MNC). Despite their potential, MNCs are vulnerable to adversarial inputs, which can lead to outages, poor decision-making, and security violations, among other issues.
@@
To address these challenges, we introduce PANTS, a practical white-box framework that uniquely integrates AML techniques with Satisfiability Modulo Theories (SMT) solvers to generate adversarial inputs for MNCs. We also embed PANTS into an iterative adversarial training process that enhances the robustness of MNCs against adversarial inputs. PANTS is 70% and 2x more likely in median to find adversarial inputs against target MNCs compared to state-of-the-art baselines, namely Amoeba and BAP. PANTS improves the robustness of the target MNCs by 52.7% (even against attackers outside of what is considered during robustification) without sacrificing their accuracy.
@@ 1 Introduction
Managing networks involves tasks such as performance routing, load balancing, and intrusion detection and is extremely challenging due to the highly diverse, heavy-tailed, and unpredictable inputs, i.e., network traffic. ML-based network-traffic classification (MNCs) provides a compelling alternative for handling these complex tasks more efficiently [26, 48, 18, 47, 32, 17, 30, 29, 61]. However, like other ML-based solutions [50, 23], MNCs are vulnerable to adversarial inputs, that is, inputs meticulously perturbed to deceive the MNC. Attackers can exploit this to compromise critical network infrastructure by subtly perturbing traffic characteristics like packet sizes and inter-arrival times, leading to security breaches or downtime. Beyond being dangerous, these attacks are particularly practical in the networking domain, where multiple entities have access to network traffic.
@@
What is critically lacking is a systematic framework for network operators and MNC designers to enhance the robustness of MNCs against attackers. This involves identifying adversarial inputs, meaning perturbations on original packet sequences that deceive the target MNC into misclassifying them. Operators can then use these inputs to: (i) refine and optimize the MNC by exploring different models, architectures, or features; (ii) assess the attack surface of their MNCs, understand the potential consequences and plan accordingly; and (iii) use them to strengthen MNCs against realistic threat models through re-training or fine-tuning. Critically, operators will not use or take the framework seriously if the generated adversarial inputs are (i) not realizable, meaning they do not fall within a realistic attacker’s capabilities, or (ii) do not retain the semantics of the original sequence, meaning the ground-truth label of the perturbed inputs has changed. More importantly, we show that adversarial inputs that are non-realizable or non-semantics-preserving can hurt the model’s accuracy when used in training.
@@
While there are numerous tools for generating adversarial samples in the Adversarial Machine Learning (AML) literature [24, 37, 14, 12, 56] they are not directly applicable to MNCs. First, AML tools are built for systems that are end-to-end differentiable, which makes them unsuitable for typical MNCs that include a feature engineering module to transform packet sequences into feature vectors before they are fed to the ML model. Although AML tools can generate adversarial feature vectors, there is no guarantee these vectors can be reverse-engineered into valid packet sequences. Even if a valid packet sequence exists, the perturbations needed to convert the original sequence into an adversarial one may exceed what a realistic attacker could achieve. Second, it is uncertain whether AML-generated samples are preserving the semantics of the original packet sequence while deceiving the MNC. Although AML techniques often incorporate a perturbation budget (expressed as a norm), this concept does not translate well to networking traffic. For example, perturbing a VoIP flow (packet sequence) to resemble web browsing by delaying packets might deceive a traffic classifier but degrades the original flow’s performance, making the call unusable. Critically, adversarially training an MNC with non-realizable or non-semantics-preserving samples (i.e., directly using what AML generates) will degrade the MNC’s accuracy without significantly improving robustness as we show in §5.2.
@@
To address these shortcomings, this paper presents PANTS, a framework to generate Practical Adversarial Network Traffic Samples that are realizable for various threat models and semantic-preserving. PANTS also enhances the target MNC through a novel iterative augmentation training process that leverages these samples. PANTS integrates traditional methods for generating adversarial samples (AML), such as Projected Gradient Descent (PGD) [37] or Zeroth-order Optimization (ZOO) [14], with formal methods, concretely a Satisfiability Modulo Theories (SMT) solver. The AML perturbs the original inputs (packet sequences or features) in a manner that maximizes their distance (e.g., loss for MLP) from the original decision boundaries with a consistently high success rate. The SMT solver finds packet sequences that do not diverge too much from the AML’s output but are realizable, consistent with the threat model and preserve the semantics of the original packet sequence. To achieve this, we encode the feature engineering, the threat model, and other networking constraints needed for semantic preservation into logical formulas. Among other optimizations, PANTS iteratively adds and removes constraints that come from the AML, prioritizing those that have a substantial effect on confusing the classifier. Finally, we integrate this generative process into an interactive training process that enhances the target MNC.
We evaluated PANTS against three distinct ML-powered network classifiers, each implemented with different models and processing pipelines. PANTS finds adversarial, realizable and semantic-preserving samples with 35.31% success rate in median, that is 70% and 2x higher than two SOTA methods. Despite being a white-box approach, PANTS works well with gradient (e.g., transformer encoder) and non-gradient (e.g., random forest) models and various processing pipelines, including those containing non-differentiable components (e.g., feature engineering). Importantly, PANTS enhances MNCs’s robustness against both white-box and black-box attackers, even when their capabilities differ from or exceed those considered during the robustification process.
We will open source PANTS and our evaluation materials to facilitate further research and MNCs robustness (cf. Open Science).
@@ 2 Motivation & Limitations of Existing Work
Figure 1: MNCs are vulnerable to adversarial inputs (circles). Traditional adversarial training sacrifices accuracy to reduce vulnerability (star). PANTS iterative adversarial-training reduces the vulnerability without hurting accuracy (cross).
@@ 2.1 Motivating use case
Consider a network operator that uses an ML-based traffic classifier to distinguish real-time applications (e.g., online gaming). The operator uses this information to optimize routing: real-time applications use fast but low-bandwidth network paths. This is one of the multiple network management tasks for which network operators could use an ML-based traffic classifier (built with Multilayer Perceptron-MLP, Random Forest, Transformers [34, 10, 26, 48, 18]) and benefit from its ability to learn complex patterns directly from data. While highly useful, this MNC exposes the network to various exploits. Two representative examples include attackers controlling (i) the end hosts of the communication or (ii) an in-network component. In the former case, a malicious (or simply “rational”) application developer could manipulate the traffic generated by her application to be misclassified as gaming by the network’s MNC, hence gaining an unfair performance advantage over competitors. This could be achieved by subtly altering the size of some packets or delaying specific packets without significantly impacting the overall performance of her application’s traffic. In the latter case, a malicious or compromised router in-path can increase the volume of traffic routed through the fast but low-bandwidth path by subtly altering the non-real-time traffic to cause the MNC to misclassify it as real-time. These perturbations would result in congestion in the low-bandwidth path, leading to delays or even downtime for actual real-time traffic, which is vital for the network’s profitability.
ML network classifiers are vulnerable to adversarial inputs. Fig. 1 illustrates the likelihood of an attacker controlling end hosts (e.g., a malicious app developer) to successfully manipulate their traffic to cause an MNC to misclassify it (predict a favorable for the attacker label). The results show that, despite the high accuracy of the models during testing, the attacker can deceive classifiers implemented with MLP, CNN, and a transformer with success rates of 81.12%, 94.25% and 99.80%, respectively. While we identified these manipulations (adversarial inputs), hence the vulnerability of the MNCs using PANTS, which is a white-box approach with complete access to the MNC’s implementation, this does not diminish the significance of the risk they pose. First, security through obscurity is never a sound strategy, and network operators should be prepared for (or at least aware of) the worst-case scenario where an attacker has complete knowledge of their system. Second, black-box approaches [35] are also able to find examples that fool MNCs, albeit with less likelihood. That is to show that adversarial examples are not random bugs but stem from learning from non-robust features, i.e., relying on unrelated co-occurrences in data instead of true correlations [28].
@@ 2.2 Requirements
Adversariness & Semantic preservation .  The framework needs to generate inputs that are adversarial, meaning that they are capable of misleading MNCs into producing an incorrect classification. To ensure that the classification is indeed incorrect, these inputs should be created by meticulously applying perturbations to the original labeled samples without altering their semantics. In other words, each adversarial input should still represent the same underlying meaning or class as the original, which is described as a set of constraints on the adversarial input. The closest notion of semantics in the image domain is the perturbation budget, which sets a maximum allowable distance between the original and perturbed image [37]. However, such a metric doesn’t translate directly to the networking domain, where there are dependencies across packets or features of each input, which perturbations should respect. We assume that semantics can be expressed as constraints on the variables of the input (not only as a distance between the adversarial and the original), and the framework should provide the means to respect them.
@@ 2.3 Limitations of existing work
Synthetic data generation Networking-specific generative models, such as NetShare [62] and NetDiffusion [31], have demonstrated strong potential in generating synthetic traffic. However, synthetic traffic is not adversarial—i.e., it should not mislead the downstream applications such as an MNC. In fact, such works aim to generate samples that match the original distribution (per the fidelity definition), whereas adversarial samples’ distribution is typically different from the training distribution [12]. This distinction is critical, as training with adversarial samples has been shown to produce more robust models compared to merely increasing the size of the training dataset [37]. As a result, NetShare [62] is 92%(99%, 92%) less effective in robustifying RF (TF, CNN) compared to PANTS and cannot robustify MLP, unlike PANTS, as we show in Fig. 9 and Fig. 14. These works may also face a semantics preservation issue: synthetically generated traces from original malicious traces (e.g., shrew attacks) might not achieve the malicious intent (e.g., cause an outage); hence might not be truly malicious.
Black-box RL-based techniques While useful for aspiring attackers, black-box RL-based approaches [22] for generating adversarial examples against MNCs are incapable of meeting our requirements. Amoeba [35], for example, trains an agent to perturb each packet to maximize the likelihood that the altered flow will deceive the MNC. Amoeba fulfills the generality requirement as it is black-box, and can fulfill realizability if trained with the right actions. Still, Amoeba’s outputs are not always semantic-preserving and adversarial because there is no way to guarantee those during inference. Also, as we show in our evaluation, Amoeba’s generated adversarial inputs are less effective in robustifying MNCs during adversarial training, possibly because it (over)exploits a vulnerability rather than trying to find new ones. Ultimately, this strategy benefits attackers but not operators seeking to enhance the MNC.
White-box gradient-based techniques Excluding AML, which is discussed in detail in §3.3, BAP [41] stands out as a notable white-box approach for adversarial generation, particularly due to its focus on networking inputs. BAP adversarially trains a generator to create perturbation on every single packet of a packet sequence (e.g., delay, add dummy bytes) such that the distance (e.g., loss) between the predicted label and the ground truth is maximized. Because BAP uses gradient for maximizing the distance, it does not work for non-end-to-end differentiable MNCs, hence failing the generality requirement. Further, by perturbing each packet independently, BAP cannot enforce semantics, which can be flow-wide. We compare against BAP in §5.
@@ 3 Overview
Figure 2: Overview of PANTS workflow. PANTS generates adversarial inputs that are also used to iteratively train the target MNC. PANTS receives the implementation of a MNC together with a training dataset and a couple of rules that constrain the generated inputs. At its core, PANTS features an AML component that collaborates with an SMT solver.
@@ 3.1 Problem formulation
We define a threat model as a set of perturbations, $\delta$ ($\delta:P\rightarrow P$) on the original packet sequences, $p_{o}$, to generate adversarial packet sequences $p_{a}$, ($p_{a}=\delta(p_{o})$). $\delta(p_{o})$ is adversarial if $f(\phi(\delta(p_{o})))\neq f(\phi(p_{o}))$, meaning that the perturbed flow will be misclassified by $f$ and $\delta(p_{o})$ preserves the semantics of $p_{o}$. Semantics are flow-level constraints on $x_{o}$ i.e., constraints that can be dependent on the whole packet sequence, not just each packet independently.
Our goal is to design a framework that generates packet sequences $p_{a}$ from $p_{o}$ such that $p_{a}$ are adversarial to $f$, semantic-preserving to $p_{o}$, and are within the capabilities of a threat model. The framework requires white-box access to $f$, hence is suitable for clouds and ISPs who develop tailored MNC solutions in-house to eliminate the need to share sensitive information externally (e.g., AT&T [5], Azure [45]) or third parties developing MNCs (e.g., Cisco [8], Juniper Networks [6]).
Figure 3: Given a network flow p, the perturbation $\delta$ is applied to generate an adversarial flow $\delta(p)$, which causes the ML model to return a wrong output.
@@ 3.2 The promise of Adversarial ML
Figure 4: PANTS combines AML with an SMT solver to generate adversarial and realizable samples (flows), which are used to assess and enhance robustness.
3.3 Adversarial ML limitations
@@
However, in the networking domain, the traffic is often processed by the feature engineering module, which generates feature vectors that are fed to the ML model. While gradient information can still guide the modification of each feature to increase loss, it is unclear how to map modifications to the feature vector into the corresponding packet sequence as the feature engineering is non-differential and non-invertible. In fact, there is no guarantee that a perturbed feature vector would have a corresponding valid packet sequence as AML changes each feature independently, making contradicting decisions. For example, suppose an ML-powered intrusion detection system whose feature set includes both minimum and maximum packet inter-arrival times ($min\_iats$, $max\_iats$). Since AML does not understand the two features are dependent, in other words, that $min\_iat\leq max\_iats$ must always hold, it is possible for AML to decrease the value of $max\_iats$ and increase $min\_iats$ such that $min\_iats>max\_iats$. In this case, while the generated feature vector might be adversarial in that it might be misclassified by the intrusion detection system, it is not a realizable sample, i.e., there is no packet sequence that would ever result in this feature set. In short, using the perturbed feature vectors is like defending against an impossible attack. While one can try to use a mask [44] to only allow AML to change a subset of the features that are not dependent, that would reduce the inputs that AML can find. Indeed, while AML can run with as few as a single feature there is no guarantee that the generated sample would be adversarial.
@@
Even if the target model does not use (non-differentiable) feature engineering or if one somehow finds a packet sequence that is consistent with the AML adversarial feature vector, the packet sequence might still fall outside the capabilities of the threat model or violate any network constraints, or break the semantics of the original flow. For example, suppose the threat model only allows injecting dummy packets (but not removing), and flow duration $f_{1}$ and packets per second $f_{2}$ are two features extracted by the feature engineering module. The number of packets can be derived from #$\#packets=f_{1}/f_{2}$. If #$\#packets$ derived from the adversarial features is less than the number of packets in the original flow, even though one can find an adversarial packet sequence consistent with the adversarial features, it still falls outside the capabilities of the threat model. Similarly, for a deep NN that is directly accepting the packet sequence as input, the AML-generated adversarial sequence could change the interarrivals in a way that is impossible for any congestion control algorithm or in a way that makes the corresponding application unusable.
3.4 PANTS: Adversarial ML for networking